OpenStudy (rachie19):

2log(4)3x+log(4) 5x = condense into a single logarithm :)

OpenStudy (jango_in_dtown):

@rachie19

OpenStudy (jango_in_dtown):

alog b=log b^a

OpenStudy (jango_in_dtown):

use this property and tell me what you can say about the first term

OpenStudy (rachie19):

Log43x^2

OpenStudy (michele_laino):

hint: for example, according to the rule of @jango_IN_DTOWN we can write: \[\huge 2{\log _4}\left( {3x} \right) = {\log _4}{\left( {3x} \right)^2}\] furthermore, please apply the subsequent rule: \[\huge {\log _4}A + {\log _4}B = {\log _4}\left( {A \cdot B} \right)\]

OpenStudy (rachie19):

Ok thanks

OpenStudy (rachie19):

So log15x^2. Or am I wrong?

OpenStudy (michele_laino):

we can write this: \[\Large \begin{gathered} 2{\log _4}\left( {3x} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ {=\log _4}{\left( {3x} \right)^2} + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left( {9{x^2}} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left\{ {\left( {9{x^2}} \right) \cdot \left( {5x} \right)} \right\} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please continue

OpenStudy (michele_laino):

what is: \(\Large 9x^2 \cdot 5x=...?\)

OpenStudy (rachie19):

45x*^2

OpenStudy (rachie19):

Log(4)45x^2

OpenStudy (rachie19):

45xcube

OpenStudy (michele_laino):

that's right! we get: \[\huge \begin{gathered} 2{\log _4}\left( {3x} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ {\log _4}{\left( {3x} \right)^2} + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left( {9{x^2}} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left\{ {\left( {9{x^2}} \right) \cdot \left( {5x} \right)} \right\} = {\log _4}\left( {45{x^3}} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

\[\huge \begin{gathered} 2{\log _4}\left( {3x} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ {\log _4}{\left( {3x} \right)^2} + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left( {9{x^2}} \right) + {\log _4}\left( {5x} \right) = \hfill \\ \hfill \\ = {\log _4}\left\{ {\left( {9{x^2}} \right) \cdot \left( {5x} \right)} \right\} = \hfill \\ \hfill \\ = {\log _4}\left( {45{x^3}} \right) \hfill \\ \end{gathered} \]